OpenStudy (anonymous):

what is the sum of all the even numbers between 88 and 0 including 88 and 0

OpenStudy (anonymous):

Okay, so \[ a_n = 0+2+4+\ldots +88 \]Let \[ b_n = \frac 12 a_n = 0+1+2+\ldots +44 \] Well \[ \sum a_n = \sum 2\times b_n = 2\sum b_n \] Such is one way to look at it.

OpenStudy (anonymous):

Likely they want you to use arithmetic sum formula... it is something like \[ S_n = n\left(\frac{a_1+a_n}{2}\right) \]

OpenStudy (anonymous):

For this formula, \(n\) is the number of terms \(45\), \(a_1\) is the first term \(0\) and \(a_n\) is the last term \(88\).

OpenStudy (anonymous):

So \[ S_{45} = 45\left(\frac{0+88}{2}\right) \]

OpenStudy (anonymous):

So that means that the integer sum of all even numbers between 88 and 0 including those is 44 * 45 which is 1980.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sum+0+to+44+of+2n Yeah, that is consistent with wolfram.